# Bayesian Quantum Amplitude Estimation

### v1

@@ III.2 Integration methods
We use Sequential Monte Carlo (SMC), a highly parallelizable method that combines a grid and MCMC to address the shortcomings of each [47, 48, 49, 50, 30, 30, 51]. The idea is to rely on a grid, but to refresh the point locations when necessary. Apart from being adaptive and efficient, this method has the added benefit of computing the model evidence at barely any extra cost. This is a valuable tool for model comparison and selection, and is not directly provided by MCMC.
Details and pseudo-code for these methods are given in appendix E.
III.3 Experimental design
As mentioned in section III, a Bayesian distribution can be used to choose the best measurements to be performed. More specifically, we can use it to evaluate the utility of any control or sequence thereof, and thus optimize it. In this work, we choose to minimize the expected variance. The target function is called a utility function. Other utility functions may be considered, such as the expected information gain [52].
Let us consider a prospective control $m$ whose benefit we want to assess. That is, we want to forecast how executing the associated circuit would improve our knowledge. We may be at a any stage of the inference process, and denote the current distribution $\mathbf{P}(\theta)$, which may consider previous data.
Being able to calculate the utility of a potential $m$, we can then optimize over a set of $M$ possible values of $m$ to choose the best option. Let us consider the case of figure 2, where we optimize from the outset the full sequence of controls based on the prior. If one intends to do $N>1$ experiments, as is usually the case, the control to be optimized is a sequence of $N$ values for $m$, and ”outcomes” are binary words of length $N$.
The cost of evaluating the utility for a single sequence of controls $\vec{m}$ grows exponentially with $N$, since there are $2^{N}$ possible outcomes whose contributions must be calculated. Let $C$ represent the cost of a full iteration of the inference algorithm without optimization, consisting essentially of a Bayesian update. The cost of calculating the utility contribution of each outcome for each control is larger than $C$. Denoting by $M$ the number of possibilities for $m$ (integers), there are $M^{N}$ possible sequences $\vec{m}$, for each of which $2^{N}$ contributions must be calculated. As such, this optimization strategy will increase the algorithm’s cost from $N*C$ to over $(M^{N}*2^{N}+N)C$. Clearly, the complexity of this approach is not favourable.
This is alleviated by a unitary look-ahead, adopting a locally optimal, or greedy, strategy. The adaptive reformulation greatly simplifies the task, reducing the cost of evaluating the utility for a sequence of $N$ experiments to linear in $N$, $(M+1)*N*C$.
A greedy approach is not guaranteed to be globally optimal, but has been shown to perform well [28]. Between the globally optimal and greedy strategies, there are multiple intermediate possibilities, where we consider look-aheads of length $L$. The cost increases exponential in the look-ahead size; the expressions we derived are special cases for $L=1$ and $L=N$.
The choice of a look-ahead size is ultimately a case-dependent negotiation between costs and benefits. The smaller the granularity, the greater the classical optimization cost, as the possible outcomes branch out exponentially; the greater the parallelism in the quantum component; and the smaller the information usage, as the knowledge used to choose the controls for the second to last experiments in a block is outdated the moment the first one is performed, and increasingly so as we advance in the block. In other words, a look-ahead of length $N$ will not make use of any of the $N$ data to inform any of the $N$ experiments (as they have been pre-determined).
Even with this simplification where the cost of one utility evaluation does not scale exponentially, the cost of optimizing it over possible choices does. To curb this, we employ problem tailored heuristics to cut down the processing costs while retaining the ability to negotiate cost to benefit trade-offs. We do this by optimizing within an exponentially expanding window that grows adaptively to suit each particular execution. We choose the utility evaluations so that the cost does not depend on the window’s width. Details are left for appendix A.
Some cost-to-benefit relations can be negotiated further. For instance, one may lower the representation cost on a continuum. In particular, the updates for the utility are not as critical as those of the main inference process. They may misguide effectiveness, but should not affect formal rigor. Additionally, one may perform multiple shot measurements for each optimized control, despite the greedy optimization. This further lowers the optimization costs. While this is not necessarily close to the optimum, good performance is observed in practice (section V).
III.4 Annealed Bayesian amplitude estimation
Until now, namely in section III.3, we considered a Bayesian inference algorithm that greedily minimizes the variance. In this section, we propose a different approach, borrowing from annealed importance sampling [50, 30], a well-known algorithm in the statistics literature.
The successive powers of the distribution form a sequence of distributions that can be sampled from using sequential Monte Carlo. Note that previously, the sequence of distributions was given by cumulative datasets of increasing size. The SMC algorithm can be applied to any sequence of distributions. Refer to appendix E for details.
IV Methods
IV.1 Processing and benchmarking
For a graphical depiction of this assessment, we will represent the root mean squared error (RMSE) as a function of the number of queries $N_{q}$. Often, tests are performed for a fixed amplitude [21, 20], which sometimes is chosen to take a specific value, such as $0.5$. This is a very particular case, which the original QAE algorithm solves exactly with just 2 auxiliary qubits. Such choices may lead to misguided assessments and overfitting, especially in statistics or machine learning based strategies.
We use an analogous expression to obtain a normalized version of the average standard deviation. This allows for a more thorough performance assessment, reveals behavior irregularities in some algorithms, and allows for universal tuning of the algorithm hyperparameters.
Note that the cost of an algorithm is calculated as the total number of queries used throughout its execution. These queries may be split among multiple circuits. This comprehensive definition of cost allows for a direct comparison with the limits of quantum metrology (section II.5), which is not concerned with the specifics of the resource allocation.
This ends the discussion on the quantities to be plotted. We now discuss how to plot them. Firstly, we will use a double logarithmic scale. This represents the limits in equations (16) and (17) as straight lines with $-0.5$ and $-1$ slopes respectively, facilitating visual assessments.
Secondly, we customize the y intercept to the datasets to facilitate a visual analysis. Details are provided in appendix C. Note that the vertical scale offset should still be heeded: it is a relevant cost metric in practice, depending on the precision regime. It may be undesirable to have algorithms requiring a higher quantum resource count from the outset, even if the scaling is Heisenberg-limited. Eventually, the performance of such an algorithm will beat that of another with a smaller offset but a less favorable evolution pace, but whether this happens depends on the target resolution.
Another challenge arises from the use of adaptivity. In quantum metrology, the performance metrics tend to respect average results, due to statistical noise. These averages are straightforward in the case of deterministic algorithms, but not with adaptive ones.
Due to a mismatch in x-coordinates, brute force averages would require discarding most data points and using a disproportional amount of executions, growing exponentially with the maximum query number. To avoid this, we find and employ a good approximation that allows us to use all data points while still depicting the statistic reliably.
We analyzed multiple possible strategies, and found the best performing one to be binning the points according to their x-coordinates and averaging both their coordinates independently.
Results for mock data are displayed in figure 5. These data were generated to reproduce ideal Heisenberg-limited behavior. By means of the aforementioned strategy, the same-colored points contained in each bin are summarized by a single point: the star markers. Ideal processing would have these markers lying unbiasedly on the dashed line. Details can be found in appendix B.
Figure 5: Results of the best performing strategy to process adaptive data (independent $x$/$y$ averaging).
IV.2 Quantum simulation
We want to numerically simulate BAE, other hybrid approaches to QAE, and the canonical QAE algorithm (to be used as a reference). For testing purposes, the quantum parts of each of these algorithms can be simulated efficiently using analytical calculations and multinomial sampling. This does not mean that the circuits can be efficiently simulated by a classical computer, as generating these data requires knowing the value of $a$. Refer to appendix D for more details.
In addition to the ideal behavior, we want to observe the behavior of the algorithms under the influence of extrinsic noise (i.e. not shot noise). We thus augment equation 9 with an extra parameter $T$ a coherence time -, and assume an exponential decay affecting the basis states equally. The factor that is independent of $\theta$ assures this symmetry (and proper normalization). The result is expression 23.
$P(\psi_{1}\mid m)=e^{-\nicefrac{{m}}{{T}}}\sin^{2}((2m+1)\theta)+\frac{1-e^{-\nicefrac{{m}}{{T}}}}{2}.$ (23)
For practicality, we define a time unit as the time taken by one application of the Grover operator. $T$ can then be expressed using these units. Having a closed expression given by equation 23, sampling noise can then be introduced as usual.
V Results
This section graphically presents the results of simulating QAE algorithms as described in section IV. Shot noise is present in all tests; in subsection V.1 no other sources of noise are considered, whereas in section V.2 a finite coherence time is imposed. The code and datasets are publicly available on GitHub [55].
V.1 Ideal simulations
Figure 6 shows the numerical simulation results for BAE. We can see that the evolution of the error is parallel to the Heisenberg limit.
Figure 6: Performance of BAE in ideal conditions (no extrinsic noise).
Figure 7 shows the results of employing the cost sparing measures described in the end of section III.3 . The offset is only slightly worse, and we still observe Heisenberg-limited estimation. The runtime for the classical processing was cut by a factor of 5.
Figure 7: Performance of BAE in ideal conditions (no extrinsic noise) with cost sparing measures.
We also test the alternative version of BAE in which, instead of locally minimizing the variance, we minimize the distance to a target effective sample size [30]. We call this version annealed BAE. Figure 8 shows the results, with the estimation still parallel to the Heisenberg limit.
Figure 8: Performance of annealed BAE in ideal conditions (no extrinsic noise) with cost sparing measures.
Finally, we test other algorithms proposed in the literature and compare them with BAE. The results are presented in figure 9. We can see that BAE has competitive performance in terms of both complexity and offset.
Figure 9: Ideal simulation results for BAE and a selection of other QAE algorithms. Acronyms describe algorithms revised in section II.6. Methodology and units are discussed in section IV.
V.2 Simulations with decoherence
To demonstrate BAE’s potential for estimation in the presence of noise, and test its and others’ behavior in the presence of noise, we run the same simulations considering a depolarizing channel with a finite coherence time. For these tests, we use random coherence times $T_{c}\in[2000,5000[$, with the time units described in section IV.2. BAE uses $500$ shots to estimate the coherence time.
The result is shown in figure 10. Our algorithm performs reliably even for higher numbers of queries, corresponding to longer executions and higher precision, whereas others with similar costs come to a stall or erratic behavior due to the noise. Among the other algorithms, the ones that don’t show this trend are those which have higher cost offsets: FAE, SAE and MLAE (LIS). This is because they’re using a higher than necessary absolute amount of quantum resources, which inflate the quantum cost necessary to achieve a fixed error. In other words, the total number of queries is not directly related to the impact of decoherence, and is not the sole predictor of unstable behavior in the presence of noise.
Figure 10: Noisy simulation results for BAE and a selection of other QAE algorithms (finite coherence time). Acronyms describe algorithms revised in section II.6. Methodology and units are discussed in section IV.
VI Conclusions and future work
We propose a method that is capable of attaining Heisenberg-limited estimation; is highly customizable and capable of negotiating trade-offs across the multiple costs involved; is parallelizable and scalable; and is resilient to noise. Numerical simulations show that our algorithm achieves Heisenberg-limited estimation and has a robust performance as compared to state-of-the-art algorithms, both in the presence and in the absence of noise.
In particular, it is capable of characterizing noise and self-adapting accordingly. This makes it better suited to faulty quantum devices than algorithms that fail to account for extrinsic noise sources. While noise can still slow down the learning rate, properly treating noise can minimize this slowdown while safeguarding correctness.
An interesting direction for future work is to test the algorithm using more complex noise models. Such models are expected to benefit the most from the robustness of the presented methods, namely the efficient multi-dimensional sampling. This improvement is a key step towards another target for future work: performing quantum amplitude estimation on real quantum devices, rather than relying on numerical simulations.
Other possible research directions include modifications to specific parts of the BAE algorithm, namely the control optimization routine or the numerical representation, to further improve the performance or reduce classical costs.
Acknowledgements.
The authors thank Ernesto Galvão, Antonio Molero and Bruna Salgado for their support and helpful comments. This work is financed by National Funds through the Portuguese funding agency, FCT - Fundação para a Ciência e a Tecnologia, within project LA/P/0063/2020. A.R. also acknowledges support from FCT under PhD grant 2022.12332.BD.
Appendix A Experimental design
As mentioned in section III, a Bayesian approach allows for estimating how informative a potential experiment is given available knowledge - that is, its utility. It does so by considering the possible scenarios that could result: how likely they are, and how useful they would be. This ability can be used to optimize the control $m$,by choosing the one with the highest expected utility.
Choosing the optimal control can then be done in a window. The simplest method would be brute force optimization over all possible controls up to a high maximum value. As mentioned in section III, the utility of each can be calculated in a look-ahead fashion, after which the one with the maximal value is chosen. The utility function could be the negative variance, the expected information gain, or other.
With this in mind, we search for the optimal control in a dynamic window that follows the exponential increase. We periodically multiply the maximum of the window by a factor. This periodically could be fixed beforehand. However, the inference algorithm is probabilistic, and hard to pre-tune in such a rigid way. In some cases, the learning will be faster than average, in which case the pre-determined schedule will slow down the learning. In less lucky runs, it may be slower, and computational resources will be unnecessarily used to evaluate the higher controls.
We obtained better results using an adaptive criterion where the expansion schedule is tailored to each run. The key idea is to expand the search window when its maximum control has been chosen and used for a measurement. As the optimal probing time tends to be exponential with the acquired information, this signals that the algorithm has captured enough information that it would benefit from a window expansion, and thus we augment the maximum.
Figure 11: Grid expansion strategy for dynamic experimental design. Green labels represent algorithm hyperparameters to be chosen by the user.
Figure 11 illustrates one occurrence of this process. In the represented case, an expansion would be triggered in the following iteration for e.g. $R=2$, $T=1$; or more generally, for $R\geq 2$, $T>1$, and one of the highest $R$ subsampled values for $m$ had already been picked for $T-1$ iterations (not necessarily consecutive).
We note that these changes reduce the optimization cost, but its exponential progression still holds, as the window width itself is expected to grow exponentially. As the search range increases, we may conjecture that the importance of choosing the exact optimum (as compared to e.g. the second best) decreases. In the beginning we have only a few options to choose from, and the trade-off between sharpness and ambiguity is particularly strong due to the absence of insight. However, once we have acquired more information and have more options to choose from, we expect the specific choice to be less critical.
With this in mind, we increase the sparseness of the searching grid as its span increases, i.e. decrease the coverage of the dynamic window as it expands. This change accomplishes a constant cost per iteration: the optimization cost increases linearly - rather than exponentially - with the number of iterations/queries. This is crucial for the scalability of the algorithm.
As mentioned in section III, we opt for a greedy strategy (unitary look-ahead). This has two advantages: leveraging all available information and minimizing the optimization cost, which would increase exponentially with the size of the look-ahead.
Pseudo-code for the discussed strategies is presented in algorithm 2. Specifics of the utility calculations can be found in appendix E.
Algorithm 2 Algorithm for optimizing an experimental control.
1: Parameters: missingNevals, $k_{0}$, $R$, $T$, $k_{\texttt{exp}}$, utility_fun
2:d $\triangleright$ Probability distribution
3:ctrl $\triangleright$ Optimized experimental control.
4:function optimize_control(d)
5:     global $c_{\texttt{min}}$, $c_{\texttt{max}}$
6:     if first call then
7:         $c_{\texttt{min}}\leftarrow$00
8:         $c_{\texttt{max}}\leftarrow$ $k_{0}*\texttt{Nevals}$
9:     else if control among $R$ highest $T$ times then
10:         $c_{\texttt{min}}\leftarrow$$c_{\texttt{max}}$
11:         $c_{\texttt{max}}\leftarrow$$2*c_{\texttt{max}}$
12:     end if
@@ Appendix A Experimental design
[…]
16:     return $\texttt{ctrl}_{\texttt{opt}}$
17:end function
For the coherence time estimation, the ’steepest’ data are associated with short evolution times, rather than long ones [22]. As such, we consider them in descending order when performing the Bayesian updates. Adaptive experimental design could also be applied to noise estimation, but we choose to use pre-determined uniformly spaced evolution times - between 0 and a provided maximum value for the coherence time. This maximum value is the only information that BAE requires to learn the coherence time, as a finite prior must be established.
Appendix B Processing adaptive data
To benchmark QAE algorithms, it is necessary to take averages over several runs for each data point. These averages are usually taken over algorithm executions, which is straightforward to do. Since these executions always use the same experiment controls and thus the same sequence of query numbers, we simply take the average of all the estimation errors for each query number. For any query number, we are sure to have as many summands as there are runs.
Once adaptivity is introduced, this is no longer as simple: the query numbers are bound to vary between runs. What is more, they tend to increase exponentially, which means that the data will become exponentially sparse as the iteration number grows. Owing to this, it would be extremely difficult to obtain averages by brute force. That is, we could run many simulations and hope that for many distinct and spaced out numbers of queries, there exist coincidental occurrences in a significant proportion of these runs. We could then represent the data points corresponding to these lucky query numbers for which we have sufficient information. However, this is extremely inefficient, especially for large iteration numbers; note that in the algorithms we have seen, the number of queries tends to grow exponentially with the iterations.
Can we do better? Ideally, we would make use of all the data, condensing it into a handful of representative data points. The aim is for the resulting graph to relay the underlying tendency correctly and intelligibly.
    $x\sim\text{unif}_{\log}([x_{\text{min}},x_{\text{max}}])$ (24)
  • Sample auxiliary variables $z$ depending on $x$:
    $z\sim\mathcal{N}(\mu,\sigma(x)),$ (25)
    where
    $\sigma(x)\propto 1/x.$ (26)
  • Calculate $y$ (representing $\epsilon^{2}$) as $y=(x-\mu)^{2}$.
$\mathbb{E}(y)=\mathbb{V}(x)\propto 1/x^{2}$ (27)
A plot of data generated like so should resemble figure 12.
  1. 1.
    Independent $x$/$y$ averaging. Average their $x$ coordinates and their $y$ coordinates separately - the naive approach. After averaging, take the square root of $\bar{y}$ to get a quantity akin to the estimation error (RMSE, or in this case standard deviation). Plot $\bar{x},\sqrt{\bar{y}}$.
  2. 2.
  3. 3.
    Average slopes. Average their $x$ coordinates to get $x$. Separately, computing the (log-)slope of each point in the group relative to the fixed point, and take the average slope $\bar{m}$. Compute the image of the mean $x$ under the power function given the average slope, and the fixed point: $f_{\bar{m}}(\bar{x})$.
We further tried two strategies without prior binning:
  1. 4.
    Curve fitting.
  2. 5.
    Spline interpolation.
Figures 13 to 17 showcase the performance of each of these approaches. An ideal candidate is one who translates each group of binned data points into a single point lying on the dashed line.
a) Noisy case.
b) Noiseless case.
Figure 13: Summary data points obtained by strategy 1: independent $x$, $y$ averaging.
a) Noisy case.
b) Noiseless case.
Figure 14: Summary data points obtained by strategy 2 - average ”log-”coordinates $\log(x)$, $\log(y)$.
a) Noisy case.
b) Noiseless case.
Figure 15: Summary data points obtained by strategy 3 - average slopes.
a) Noisy case.
b) Noiseless case.
Figure 16: Summary data points obtained by strategy 4 - curve fitting.
a) Noisy case.
b) Noiseless case.
Figure 17: Summary data points obtained by strategy 5 - spline interpolation.
Independent $x$/$y$ averaging seems to yield the best results; accordingly, we adopt this strategy where necessary.
Appendix C Defining the intercept
To plot data along with the fundamental limits of section II.5, we must define the y-intercept. This is not given by these limits alone: they describe only the slope (in a log-scale), which represents the complexity.
  • Do a curve fit $Be^{m}$ with parameters $B$, $m$:
    $y=Bx^{m}\leftrightarrow\log(y)=m\log(x)+\log(B)$
This seems to produce sensible results.
Appendix D Numerical treatment of QAE
D.1 Original algorithm (QFT)
$P(\text{measuring }x\mid QPE(\phi))=\frac{\sin^{2}(K\Delta\pi)}{K^{2}\sin^{2}(\Delta\pi)},$ (28)
$\Delta=\left\lvert\phi-\frac{x}{K}\right\rvert\mod 1.$ (29)
This is an angular distance, divided by $2\pi$.
    • The exact case outcome is $x_{0}=M\theta/\pi$.
    • We would calculate $\theta$ from an outcome $x$ as $\pi x/M$.
    • $\Delta$ is defined as $\Delta=\lvert 1-\theta/\pi-x/M\rvert\mod 1$, which measures the distance between $1-\theta/\pi$ and $x/M$; or equivalently, between $\theta/\pi$ and $1-x/M$.
The second points in each of the two cases, which tell us how to calculate the unknown parameter $\theta$ from an outcome $x$, may seem problematic: in practice, we have no way of knowing which eigenphase was behind a measurement. However, as luck would have it, the object of our interest is not actually $\theta$, but rather $a=\sin^{2}(\theta)$. Due to the symmetry of the sine function about $\pi/2$ in $[0,\pi]$, $a$ can always be calculated as $a=\sin^{2}(\pi x/M)$, regardless of the underlying eigenvalue. Equivalently, it can be calculated as $a=\sin^{2}(\pi-\pi x/M)$.
$P(x\mid QAE(\theta))$ $=\frac{P(x\mid QPE(\theta/\pi))}{2}$ (30)
$+\frac{P(x\mid QPE(1-\theta/\pi))}{2}$
$P(x\mid QAE(\theta))$ $=\frac{P(x\mid QPE(\theta/\pi))}{2}$ (31)
$+\frac{P(M-x\mid QPE(\theta/\pi))}{2}.$
D.2 Maximum likelihood post processing
If we execute the original QAE algorithm for one of the few cases allowing an exact result, we will always get the same outcome. But for most cases, we will get a distribution of outcomes, according to section D.1. One must then pick one estimate for the amplitude.
D.3 Grover measurements
Similar to section D.1, measurement data for Grover measurements can be generated efficiently. In this case, the probabilities are straightforward and given by equation (9). To add shot noise, one need only draw from the Bernoulli distribution defined by this parameter.
Appendix E Sequential Monte Carlo
Pseudo-code for a Sequential Monte Carlo update is presented in algorithm 3. Note that normalization is not necessarily required after each update, as it is only required when calculating quantities (algorithm (5).
1: Parameters: N, $T$triangleright Number of particles, resampling threshold.
2:$\vec{\theta}$, $\vec{w}$, $\vec{m}$, $\vec{o}$triangleright Particle positions and weights, controls, outcomes.
3:$\vec{\theta}$’, $\vec{w}$$\triangleright$ Updated positions and weights.
@@ Appendix E Sequential Monte Carlo
[…]
@@
12:     for $i\in\{1,\,\dots,N\}$ do
13:         $\vec{w}[i]\leftarrow$ $\vec{w}[i]*\mathcal{L}(\vec{\theta}[i]\mid\texttt{o};\texttt{ctrl},N)$triangleright Bayes’ rule.
@@
[…]
16:     if ESS($\vec{w}$) $>$ $T$ then
[…]
@@
2:$\vec{\theta}$, $\vec{w}$, info $\triangleright$ Particle positions and weights, information for the kernel.
3:$\vec{\theta}$’, $\vec{w}$$\triangleright$ Updated positions and weights.
4:function resample($\vec{\theta}$, $\vec{w}$)
@@
[…]
@@
[…]
@@
1:$\vec{\theta}$, $\vec{w}$, f $\triangleright$ Particle positions and weights, and a function.
@@
3:function expectation($\vec{\theta}$, $\vec{ws}$, f)

### v5

@@ Appendix B Integration Methods
We use Sequential Monte Carlo (SMC), a highly parallelizable method that combines a grid and MCMC to address the shortcomings of each [11, 12, 31, 37, 47, 47, 10]. This type of sampler has been applied to a multitude of tasks in quantum metrology [20, 17, 57, 55, 45], although using the Liu-West filter (LWF) [31] instead of MCMC. The underlying idea is to rely on a grid, but to refresh the point locations by resampling (with e.g. LWF or MCMC) when necessary. Apart from being adaptive and efficient, this method has the added benefit of computing the model evidence at barely any extra cost. This is a valuable tool for model comparison and selection, and is not directly provided by MCMC. In the simulations realized in this work, we use MCMC rather than LWF, as it is more robust and provides more accurate expectations, even e.g. under multi-modality.
Details and pseudo-code for these methods are given in Appendix B.1.
B.1 Sequential Monte Carlo
Pseudo-code for a Sequential Monte Carlo update is presented in algorithm 2. Note that normalization is not necessarily required after each update, as it is only required when calculating quantities (algorithm 4). These algorithms are one possible solution for the unspecified functions in algorithm 1; in this case, distributions are represented by sets of points $\vec{\theta}$ (often called particles in SMC literature) and their respective weights, $\vec{w}$. For details on the hyperparameters Npart, thr, refer to Appendix J.
1: Parameters: missingNpart, thr $\triangleright$ Number of particles, resampling threshold.
@@ B.1 Sequential Monte Carlo
7:   $\texttt{$\vec{\theta}$, $\vec{w}$}\leftarrow$ bayesian_update($\vec{\theta}$, $\vec{w}$, $\vec{m}$[i], $\vec{o}$[i])
@@
12:  for $i\in\{1,\,\dots,\texttt{Npart}\}$ do
13:   $\vec{w}[i]\leftarrow$ $\vec{w}[i]*\mathcal{L}(\vec{\theta}[i]\mid\texttt{o};\texttt{m})$ $\triangleright$ Bayes’ rule.
@@
15:  $\text{ESS}\leftarrow$ $\left(\sum_{i=1}^{N}\texttt{w[i]}\right)^{2}\bigg{/}\left(\sum_{i=1}^{N}\texttt{w[i]}^{2}\right)$ $\triangleright$ Effective sample size.
16:  if ESS($\vec{w}$) >> thr then
17:   $\texttt{$\vec{\theta}$', $\vec{w}$'}\leftarrow$ $\vec{\theta}$, $\vec{w}$
@@
4:function resample($\vec{\theta}$, $\vec{w}$)
5:  $\texttt{N}\leftarrow$ length($\vec{\theta}$)
6:  $\vec{\theta}^{\prime}\leftarrow$ empty list
@@
8:   $\theta\leftarrow$$\vec{\theta}$[j] sampled with probability $\vec{w}$[j] $\triangleright$ Multinomial sampling.
9:   $\theta^{\prime}\leftarrow$ kernel($\theta$, info)
@@
12:  $\vec{w}^{\prime}\leftarrow$ array of ones with length N
@@
[…]
@@
3:function expectation($\vec{\theta}$, $\vec{w}$, f)
4:  $\texttt{S}\leftarrow$ $\sum_{i}\vec{w}$[i] $\triangleright$ Normalization factor.
5:  $\texttt{E}\leftarrow$ E + $\sum_{i}$ f($\vec{\theta}$[i]) * $\vec{w}$[i]
6:  return E/S
7:end function
[…]
2:E $\triangleright$ Average expectated value for f for $m$.
3:function average_expectation($\vec{\theta}$, $\vec{w}$, f, $m$)
4:  $\texttt{E}\leftarrow$ 0
5:  for each possible outcome o do
6:   $\texttt{P}\leftarrow$ expected_probability($\vec{\theta}$, $\vec{w}$, $m$,o)
[…]
8:  end for
9:  return E
10:end function
11:function cond_expectation($\vec{\theta}$, $\vec{w}$, f, o, $m$)
12:  $\texttt{$\vec{\theta}$', $\vec{w}$'}\leftarrow$ bayesian_update($\vec{\theta}$, $\vec{w}$, $m$, o)
13:  $\texttt{cE}\leftarrow$ expectation($\vec{\theta}$’, $\vec{w}^{\prime}$, f)
14:  return cE
15:end function
16:function expected_probability($\vec{\theta}$, $\vec{w}$, $m$,o)
17:  $\texttt{S}\leftarrow$ $\sum_{i}\vec{w}[i]$ $\triangleright$ Normalization factor.
18:  $\texttt{P}\leftarrow$ 0
19:  for $i\in\{1,\,\dots,N\}$ do
20:   $\texttt{P}\leftarrow$ P + $\vec{w}[i]*\mathcal{L}(\vec{\theta}[i]\mid\texttt{o};$m))
21:  end for
22:  return P/S
23:end function
B.2 Liu-West filter
The Liu-West filter is a resampling kernel that preserves the two first moments of the distribution [31]. Pseudo-code is presented in algorithm 6. It may perform well enough for simple problems, while being more lightweight than Markov Chain Monte Carlo.
Algorithm 6 Algorithm for a Liu-West resampling kernel.
1: Parameters: $\alpha$